OpenStudy (anonymous):

4 + m/8 = 3/4 = ? (Show Work)

OpenStudy (dumbcow):

\[4+\frac{m}{8} = \frac{3}{4}\] correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Clear fractions... multiply everything by the LCD... which would be 8.

OpenStudy (anonymous):

Let's see what you get after clearing fractions... or perhaps ask a question.

OpenStudy (anonymous):

like how?

hero (hero):

Why not just multiply 3/4 by 2/2

OpenStudy (anonymous):

\[(8)(4+\frac{m}{8})=\frac{3}{4}(8)\]Distribute on the left, mulitpy by reducing on the right.

hero (hero):

If you multiply 3/4 by 2/2, you'd have \[4 + \frac{m}{8} = \frac{6}{8}\]

hero (hero):

Then you could subtract m/8 from both sides afterwards.

OpenStudy (anonymous):

Clearing fractions is a standard technique in intermediate algebra so that the student can avoid the fears that come from dealing with fractions. You can, of course, manipulate fractions, but the student usually prefers to avoid it unless absolutely necessary.

hero (hero):

I was never a fan of standard techniques.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So which would you like, clearing fractions, or work with them... either is totally valid.

hero (hero):

That's like a 'trick question'.

OpenStudy (anonymous):

No, really... I like fractions... either is ok.

hero (hero):

You should probably ask, would you like to work with LCDs or regular algebra.

OpenStudy (anonymous):

LCD's are regular algebra. It is strategically multiplying both sides by the same number to get a manageable result. Balancing equations are at the heart of algebra.

OpenStudy (anonymous):

Standard techniques provide students who are below college algebra level with a sense of security. At the college algebra level, numerous methods should be presented so that the student becomes adaptable to various situations. Before that, it generally freaks them out.

hero (hero):

I honestly would rather avoid calling it LCDs because it tends to confuse students.

hero (hero):

You're trying to suggest that no student have problems with LCDs.

OpenStudy (anonymous):

I can type out lowest common denominator if necessary.

OpenStudy (anonymous):

No, I am just using the advice given to me by my employer for dealing with students at various levels. I see what students encounter in their courses every day, so I try to emply those techniques.

OpenStudy (anonymous):

I sense that you solve problems with a highly intuitive approach supported by solid technique. I love students like you. You advance math. But I've seen scared students too. I try to help them through.

hero (hero):

And in this case, I try to show students that you can apply simple algebra to solve problems. Yeah, I have all kinds of alternative methods I use.

OpenStudy (anonymous):

We lost our student. :( I hope that she got it.

hero (hero):

What a pity. I hope so.

hero (hero):

Don't put me in a classroom

OpenStudy (anonymous):

I hope that I haven't aggravated you Hero. I appreciate that contributions that I've seen you provide. I simply have an approach that I have been trained to provide. I am sometimes winging it, but at this level I have tutored many students and feel confident. Please don't believe that I hold any ill will.

hero (hero):

"trained to provide". So does that mean you are not allowed to show alternative methods?

hero (hero):

If I were in a classroom, I would show students the standard and alternative methods

hero (hero):

And let them choose which one they prefer

OpenStudy (anonymous):

Oh, I absolutely am, but students below college algebra level generally want to know... what is step 1... then step 2... then step 3... just like my teacher showed me or "la la la I can't hear you". It is very common at the community college level. They tune you out or lose faith if it doesn't look "just like my teacher said". At college algebra, I try to open them up. Look at it from numerous perspectives and develop a style.

OpenStudy (anonymous):

Then I don't care what their teacher says. Their are many perspectives, choose the one that works for you. Fit the teacher enough to pass the class and move beyond.

hero (hero):

Of course, I would show both methods and wouldn't forget to put "step 1" "step 2" in my presentations.

hero (hero):

Some "below college level" students like the alternative methods.

OpenStudy (anonymous):

I suspect that you appreciate the multiple perspective approach highly, but you are not in the majority. Most get scared by that.

OpenStudy (anonymous):

And I don't mean college level exactly. I mean College Algebra level. In high school it is commonly referred to as Algebra 2.

hero (hero):

I suppose Algebra II is what you teach.

OpenStudy (anonymous):

It's just a title. You need "College Algebra" to get the associate's degree and often it is all you need for a liberal arts degree.

hero (hero):

Oh, I see. To be honest, I did learn the alternative methods later on, but now I'm mad at the teachers I had for not knowing them or showing them to me. I think that students who have already mastered a certain concept could definitely benefit from alternative methods.

OpenStudy (anonymous):

I agree. I just don't want to scare them away. I want them to learn to love it, then open the door to see more.

OpenStudy (anonymous):

I think that there are completely alternative ways to learn this stuff that would work well. I'm just trying to work with the system that is in place.

OpenStudy (anonymous):

This site is awesome for finding other ways.

hero (hero):

I agree. I am pretty much an alternative methods geek. I'm always looking for alternative methods. One day, I might right a book with nothing but alternative methods in them.

OpenStudy (anonymous):

That would be a great contribution. Like I said, it is people like you who advance math rather than just learn what is already there and regurgitate. I will admit that I am generally a regurgitator that admires new things to regurgitate.

OpenStudy (anonymous):

If I ever get a totally fresh idea on my own, that satisfaction would probably put a smile on my face for years.

hero (hero):

Yeah, I'm like that too. That's why I'm more than excited to show the alternative methods.

hero (hero):

I always seem to develop new methods that spring from annoyance with the "by the book" methods. I'm always convinced that there is some other method that works that could be more straightforward or easier. The way I solve the fractions is one method. Then there's the method for finding the slope of a line. But I have plenty other alternative methods too.